OpenStudy (anonymous):

Use synthetic division to find the quotient and remainder when \[x^6-1\] is divided by \[x+1\]

mathslover (mathslover):

\[\huge{(x^3)^2-(1)^2=[{(x^3)+1}][{x^3-1}]}\]

OpenStudy (anonymous):

 \[x^6-1=(x^5)(x+1)-2\]

OpenStudy (anonymous):

What did you do @mathslover

OpenStudy (anonymous):